Given the production function:
Q = 100 + P – 0.01P2 + 2N – 0.03N2
Determine the marginal rate of technical substitution
1
Expert's answer
2012-03-15T12:49:56-0400
Marginal Rate of Technical Substitution (or Technical Rate of Substitution) is
the amount by which the quantity of one input (n) has to be reduced when one
extra unit of another input (n2) is used, so that the total output (q) remains
constant:

Dn/Dn2 = MP(n2)/MP(n) = -0.03/2= -0.015 – MRTS for the input n
for n2